Crane Company receives a $86,000, 5-year note bearing interest of 5% (paid annually) from a customer at a time when the discount rate is 6%.

Click here to view the factor table.

(For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

What is the present value of the note received by Crane? (Round answer to 2 decimal places, e.g. 25.25.)

Present value of note received

$enter the present value of the note received in dollars rounded to 2 decimal places
Answer #1

Compute the present value of the note received:

Interest [$86,000 \times 5%] = $4300 per annum

PVAF(6%,5) = 4.21236

PVF(6%,5) = 0.74726

PV = $4,300 \times 4.21236+ $86,000 \times 0.74726

= $18,113+ $64264

= $82,377